[Brahms-l] Shift report 20050128 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Jan 28 2005 - 00:03:33 EST
Started the shift with Hiro working on Si detectors.  Due mainly to an alledged short in the gate connector, we had problems which stayed no matter which modules were inserted.  When this was finally diagnosed, we believe that si detectors are OK.  Hiro will provide more information on exactly what was corrected.

Beam delivered at 18:30
Started taking data at 20:15

Changed D5 to A polarity.
Ran FS at 8 deg A1219 and MRS 90 deg A2255
Runs     tr2      tr3
12540   74k       51k
12542   57k       40k
12543   36k       25k
Beam dumped at 23:30

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jan 28 00:04:02 2005

This archive was generated by hypermail 2.1.8 : Fri Jan 28 2005 - 00:04:26 EST